the half life of a first order reaction is 100 seconds at 280K. if the temperature coefficient is 3.0 its rate constant at 290K in 1/s is???

Grade:12

1 Answers

Gundeep Singh
7 years ago
As Temperature coefficient is 3, so (rate const. at 290k)/(rate const. at 280k)=3 rate const. at 280k = ln2/100 =0.00693 /s rate const. at 290kk = 3ln2/100=0.0208 /s